OpenStudy (anonymous):

find the sum x^3-2y^3 (+)4x^3+3x^2y-2y^2+3y^3

jhonyy9 (jhonyy9):

x*3 -2y*3 +4x*3 +3x*(2)y -2y*2 +3y*3 =5x*3 +y*3 +3x*2y -2y*2= =x*2(5x+3y)+y*2(y-2)